# The 2010 Census Confidentiality Protections Failed, Here’s How and Why

### v1

@@ 8.2. Detailed Results for Nonmodal and Modal Race-Ethnicity Records by Census Block Size
Figures 3 and 4 summarize the results, by census block size, for the low quality COMCRL and higher quality $\text{CEF}_{atkr}$ data, respectively. The detailed data are shown in Appendix Tables 12, 13, and 14. The first column of both figures shows the results for all data-defined persons by census block size. As summarized in the national results discussion, except for the lowest population blocks (1 to 9 persons), there are very few discernible differences between the CEF, HDF, $\text{rHDF}_{b,t}$, $\text{rHDF}_{b}$, MDG, and PRG results. Comparing the results for nonmodal data-defined persons (middle column) with those of modal data-defined persons (last column) reveals the power in the reconstructed microdata vis-á-vis the statistical baselines. There is only one substantive analytic difference between the COMCRL and $\text{CEF}_{atkr}$ results. Because we use the CEF itself to determine a nonmodal person, the data-defined populations in the COMCRL nonmodal and modal columns are a strict subset of the data-defined population in the all persons column: only the 106,300,000 records in the top left cell of Table 3 can be used to contrast the vulnerable (nonmodal) and nonvulnerable (modal) populations, whereas all 286,700,000 data-defined COMCRL records in the CEF universe can be used for the all persons column. All data-defined persons in the CEF can be studied in the $\text{CEF}_{atkr}$ results. For both attacker models, the denominators in the nonmodal and modal putative and confirmed reidentification rates from equations (11) and (12) are the actual at-risk data-defined subpopulations (nonmodal and modal, respectively). The reidentification precision rate defined in equation (13) is unaffected by restricting the domain of COMRCL. An external attacker using either COMRCL or higher quality data would still have to make inferences directly from the reconstructed microdata or form independent estimates of the race and ethnicity of the census block to estimate the putative or confirmed reidentification rate. But the external attacker’s estimate of the precision rate would still depend only on independent field work using a labeled sample of the attacker’s putative reidentification as noted in Section 5.2.
Figure 3. Comparison of Putative Reidentification Rates, Confirmed Reidentification Rates, and Reidentification Precision Rates for COMCRL by Census Block Size
Notes: The denominator used in the first column is the COMCRL Population column in Table 12, which totals to 286,700,000. The denominator in the second column is the COMRCL population in Table 13, and the denominator in the third column is the COMRCL population in Table 14. The denominators in the second and third columns sum to 106,300,000 because only that subset of COMRCL records can be classified as modal and nonmodal from the CEF. See Table 3.
@@
Notes: The denominator used in the first column is the $\text{CEF}_{atkr}$ Population column in Table 12, which totals to 276,000,000. The denominator in the second column is the $\text{CEF}_{atkr}$ population in Table 13, and the denominator in the third column is the $\text{CEF}_{atkr}$ population in Table 14. The denominators in the second and third columns sum to 276,000,000 because all records in $\text{CEF}_{atkr}$ can be classified as modal and nonmodal from the CEF. See Table 3.
For comparing and contrasting the nonmodal and modal persons results, the interpretation does not depend on the attacker database, so we characterize them generically. In these figures, nonmodal data-defined persons are the vulnerable population. Except for blocks with populations of 1-9 persons, CEF, HDF, $\text{rHDF}_{b,t}$, $\text{rHDF}_{b}$, MDG, and PRG putatively reidentify nonmodal persons at essentially the same rates because SF1 Tables P12 and 14 contain all the information used for putative reidentifications. It is the confirmation and precision rates that reveal the contrast with the modal person subpopulation. In blocks with 1-9 persons, the HDF, $\text{rHDF}_{b,t}$, and $\text{rHDF}_{b}$ have essentially identical confirmation and precision rates, whereas the MDG baseline is nearly zero for both confirmation and precision rates, and the PRG baseline is greater than the MDG but substantially less than either $\text{rHDF}_{b,t}$ or $\text{rHDF}_{b}$. As the population in the census block increases, the confirmation and precision rates of $\text{rHDF}_{b,t}$ and $\text{rHDF}_{b}$ decline but remain high, much closer to the rates for CEF and HDF than MDG (always essentially zero) and approaching proportional guessing only for the most populous census blocks. Even though there are 125 possible nonmodal {race, ethnicity} combinations, the reidentification precision rate of the reconstructed data is greater than 50% for 23,765,000 persons living in census blocks with populations of 1-99 persons (more than $1/3$ of all nonmodal data-defined persons) compared to essentially 0 for the modal guesser. For this same group, the smallest gain in the reidentification precision rate relative to the proportional guesser is 36.8 percentage points. This enormous gain in accuracy occurs only because the nonmodal person’s data were used in the SF1 tabulations. When the vulnerable population is defined as persons who differ from the predominant characteristics in small neighborhoods, in this case nonmodal {race, ethnicity} persons in lower-population census blocks, the statistical baselines either fail completely (MDG) or fair no better than chance (PRG), but the reconstructed microdata are correct at rates that approach the correctness of the confidential data themselves as the neighborhood population decreases.
@@ 8.3. Results for the Most Vulnerable Populations
We now discuss the results for the most tightly constrained definition of the vulnerable population—nonmodal data-defined persons living in blocks with 0 solution variability whose reconstructed records are population uniques for {block, sex, agebin}. Panel A of Table 10 shows the results for nonmodal persons living in 0 solution variability blocks and Panel B shows the results for the subset of these persons who are population unique on {block, sex, agebin}. Recalling that population uniqueness on {block, sex, agebin} implies population uniquness on {block, sex, age} and that 0 solution variability implies that the record in the reconstructed data is provably the same as the record in HDF without referencing any confidential data, Table 10 establishes that the reidentification precision rates of $\text{rHDF}_{b,t}$ and HDF are essentially identical 91.2 versus 93.0 in COMCRL and 95.2 versus 95.6 in $\text{CEF}_{atkr}$. These reidentification precisions are far in excess of the modal guesser’s (2.5 at best) or the proportional guesser’s (20.2). Thus, the correct {race, ethnicity} has been exposed for more than 19 of 20 persons in the population of nonmodal, 0-$solvar$, {block, sex, agebin} uniques—essentially the same exposure as publishing the HDF itself, and only five percentage points worse than publishing the CEF itself. Statistical baselines are either almost always wrong or wrong for 4 of 5 persons in this vulnerable population. These results definitively show that the published SF1 tables result in confidentiality breaches since the high statistical precision of the reconstructed microdata is only possible because the target person’s data were used in the SF1 tabulations. If HDF was unsafe to release under the 2010 Census disclosure limitation framework, then the Summary File 1 statistics in Table 2 were also unsafe to release.
@@
Algorithm 3) (×$\times 10^{-3}$) (×$\times 10^{-3}$) (×$\times 10^{-3}$) (%) (×$\times 10^{-3}$) (×$\times 10^{-3}$) (×$\times 10^{-3}$) (%)
@@
requirements. The rows labeled $\text{rMDF}_{b,t}$ and MDF (light gray highlight) are the analogues
of $\text{rHDF}_{b,t}$ and HDF, respectively, when the 2020 Census DAS is applied to the 2010 CEF.
The rows labeled $\text{rSWAPLo}_{b,t}$ and $\text{rSWAPHi}_{b,t}$ (medium gray highlight) are the analogues
@@
of $\text{rHDF}_{b,t}$ when alternative swapping parameters are used. See Section 10. Population for
@@
attacker COMRCL is data-defined nonmodal person records in COMRCL that match CEF
@@
on the feature set {pik, block}. Population for attacker $\text{CEF}_{atkr}$ is all data-defined
nonmodal persons in CEF.
9. The 2010 Census Statistical Disclosure Limitation Framework Did Not Meet the Census Bureau’s Own Standards
@@
The results of our experiment show that it is possible to create an extremely accurate reconstruction (95% binned-age agreement, see Table 8) of the underlying HDF microdata from publicly available sources, effectively undoing any confidentiality protection resulting from releasing tabular summaries instead of releasing the same features as a microdata extract of the universe of HDF person records. There was no reconstruction solution variability for $70\%$ of populated blocks, meaning that the reconstructed records in $\text{rHDF}_{b,t}$ were the exact image of the records in HDF for the feature set {block, sex, agebin, race, ethnicity}. For these blocks, housing-unit swapping and synthetic data for the group quarters population are the only sources of uncertainty about their correctness relative to the CEF. These SDL measures do affect the accuracy of the reconstruction relative to the CEF, especially in the blocks with the smallest population sizes (74% agreement in blocks with population size of 0-9), but the accuracy is still very high overall (and could be higher for attackers who attempt to undo the swapping). When the reconstructed data are linked to an external dataset, an attacker can learn extra characteristics about members of their dataset that they previously did not know, and could not learn statistically, with very high accuracy for select publicly identifiable subgroups. The attacker’s successful inference depends on several factors. First, the more accurately the external dataset matches the linking variables {block, sex, agebin}, the more records the attacker will successfully link and the better the inference about the value of the other census responses will be. In our experiment, only about $40\%$ of the records in our low quality commercial data were correct matches to the CEF on these linking variables. Consequently, even though we could make a large number of putative matches, many were not confirmed–not because of incorrect inference on the unknown characteristics {race, ethnicity}, but because the characteristics on the external dataset were inaccurate quasi-identifiers. In our experiments using a much more accurate external database (either the $\text{CEF}_{atkr}$ or the subset of records in the commercial database that matched the CEF) the confirmation and precision rates were much higher. As the accuracy of the external data improves, the success of our attack improves, indicating that any conclusions based on the relatively low quality of commercial name and address databases in 2010 should be tempered by the improved quality of such data by 2020.
@@
Our reconstruction-abetted reidentification atack was much more accurate in successfully inferring the $\{race,ethnicity\}$ of records with a value that was different than the most common (modal) $\{race,ethnicity\}$ in the block in comparison with the baseline inferential methods of guessing the modal {race, ethnicity} and guessing the {race, ethnicity} with probability proportional to publicly released counts. It also vastly outperformed the baseline methods when making inferences about records that were unique and had nonmodal $\{race,ethnicity\}$, facts that the attacker could infer exactly in blocks with zero solution variability without consulting any confidential data. Therefore, while it is possible to make accurate inferences with these baseline methods when statistical information is used, the reconstruction-abetted reidentification attack exposes the vulnerability of the confidentiality protections by permitting accurate inferences on persons with statistically uncommon and rare characteristics. The accurate inference of the race and ethnicity of a person who is a population unique is only possible because that person’s record was available in $\text{rHDF}_{b,t}$. Such an inference is nonstatistical using the framework laid out in Section 5. Record-level data swapping marginally reduced the number of records in the $\text{rHDF}_{b,t}$ at risk for a putative match, but it had no discernible effect on the precision of nonstatistical inferences for the vast majority of records.
@@
The conclusion from this analysis is that the disclosure avoidance system applied to the 2010 Census did not meet the requirements for approval set forth by the Census Bureau’s Data Stewardship Executive Policy (DSEP) Committee for 2010 Census publications and, in light of current algorithmic and computational capabilities, it is not viable for future use. The results of the this experiment motivated the Census Bureau to change the SDL methods used for the 2020 Census [abowd2018us]. Instead of record swapping, synthetic group quarters data, and aggregation, the Bureau used differentially private mechanisms for all tabulations from the 2020 Census [Abowd20222020]. Had the Census Bureau continued to use the 2010 technology for 2020, it would have had to use the same rules for microdata as for tabular data, likely resulting in a much higher swap rate and a large amount of suppression of block-level data.
10. The 2020 Census Statistical Disclosure Limitation System Addressed the Failures of the 2010 System
@@
Our final discussion addresses whether any SDL method can simultaneously defend against reidentification attacks and other nonstatistical uses while allowing publication of data that are fit for use in the primary applications. [christ:radway:bellovin:2020, ] demonstrated that the Census Bureau’s differentially private disclosure avoidance system could provide fitness-for-use in the redistricting application comparable to swapping but with superior confidentiality protection. They also noted that “[s]wapping places a disproportionate privacy burden on minority groups,” while differential privacy protections apply to all sub-populations. These results, which we confirm in this paper, are in contrast to the erroneous conclusions drawn by [kenny:et:al:2021, ] who calculated percentage changes incorrectly for small populations and failed to distinguish between generalizable and privacy-violating inferences [jarmin:etal:2023]. [wright:irimata:2021, ] demonstrated that the 2020 Census Redistricting Data (P.L. 94-171) Summary File was fit for use in redistricting and scrutiny under Section 2 of the 1965 Voting Rights Act.
@@
On October 20, 2022, DSEP approved the final production settings for the 2020 Census disclosure avoidance system as applied to the Demographic and Housing Characteristics (DHC) File, the successor to the 2010 Summary File 1 used in this paper. DHC was produced using an extension of the algorithms described in [Abowd20222020, ] that includes the 2020 Census version of the tables used in this paper [cumings-menon:etal:2023]. The statistical summaries presented for the DSEP decision included an analysis of the effectiveness of the same reconstruction-abetted reidentification attack studied here based on the 2010 Census. The analysis compared the confidentiality protections from the original SF1 tables shown in Tables 2, 3 and 4 with the protection afforded by the differentially private disclosure avoidance system, when executed at the DHC production settings applied to the 2010 Census data, and the protection afforded by housing unit-level data swapping at rates of 5% and $50\%$, i.e., swapping 5% or half of all housing units to a new location in a different block, tract, or elsewhere in the state. See Appendix C for details.
@@
Table 11 shows the agreement rates overall and by census block size between the 2010 CEF and
@@
  • $\text{rMDF}_{b,t}$—reconstructed Microdata Detail File (MDF), [Abowd20222020], the production version output of the differentially private DAS for the DHC when applied to 2010 Census data and processed into the equivalents of the Table 2 statistics,
@@
The $\text{rMDF}_{b,t}$, MDF, and $\text{rSWAPHi}_{b,t}$ reduce the agreement with the 2010 CEF substantially, with the $\text{rMDF}_{b,t}$ and MDF reducing agreement more than $\text{rSWAPHi}_{b,t}$. The $\text{rSWAPLo}_{b,t}$ agrees with the CEF at essentially the same rate as $\text{rHDF}_{b,t}$ except for blocks with populations of 1-9 persons, where HDF, $\text{rHDF}_{b,t}$, and $\text{rHDF}_{b}$ all agree with the CEF for 74.0% of records using binned age while $\text{rSWAPLo}_{b,t}$ agrees with 94.9% of these records. The experimental swap methodology did not single-out low population blocks for relatively greater swapping.
Table 11. Selected Reconstruction Agreement Statistics with Comparisons to Output from the 2020 Census Disclosure Avoidance System and Specially Swapped Versions of the CEF Using the 2010 Census as Input by Census Block Size
@@
Data ($L-R$ in Block Population Agreement (×$\times 10^{-3}$) Agreement (%)
Algorithm 2) Size (×$\times 10^{-3}$) Exact Age Binned Age Exact Age Binned Age
@@
$\text{rSWAPLo}_{b,t}$-CEF All 308,746 144,500 281,200 46.8 91.1
$\text{rSWAPHi}_{b,t}$-CEF All 308,746 100,800 193,200 32.7 62.6
@@
$\text{rSWAPLo}_{b,t}$-CEF 1-9 8,070 3,278 7,660 40.6 94.9
$\text{rSWAPHi}_{b,t}$-CEF 1-9 8,070 1,803 4,280 22.3 53.0
@@
$\text{rSWAPLo}_{b,t}$-CEF 10-49 67,598 30,320 63,370 44.9 93.8
$\text{rSWAPHi}_{b,t}$-CEF 10-49 67,598 18,110 38,330 26.8 56.7
@@
$\text{rSWAPLo}_{b,t}$-CEF 50-99 69,073 31,190 63,300 45.2 91.6
$\text{rSWAPHi}_{b,t}$-CEF 50-99 69,073 20,200 41,180 29.2 59.6
@@
$\text{rSWAPLo}_{b,t}$-CEF 100-249 80,021 36,310 71,880 45.4 89.8
$\text{rSWAPHi}_{b,t}$-CEF 100-249 80,021 25,740 50,530 32.2 63.2
@@
$\text{rSWAPLo}_{b,t}$-CEF 250-499 42,911 20,470 38,250 47.7 89.2
$\text{rSWAPHi}_{b,t}$-CEF 250-499 42,911 15,830 29,030 36.9 67.7

### v2

@@ 7.2. Detailed results for nonmodal and modal race-ethnicity records by census block size
Figures 3 and 4 summarize the results, by census block size, for the lower quality COMRCL and higher quality $\text{CEF}_{atkr}$ data, respectively. The detailed data are shown in Appendix Tables LABEL:tab:reid_blockpop, LABEL:tab:reid_blockpop_nm, and LABEL:tab:reid_blockpop_m. The first column of both figures shows the results for all data-defined persons by census block size. As summarized in the national results discussion, except for the lowest population blocks (1 to 9 persons), there are very few discernible differences between the CEF, HDF, $\text{rHDF}_{b,t}$, $\text{rHDF}_{b}$, MDG, and PRG results. Comparing the results for nonmodal data-defined persons (middle column) with those of modal data-defined persons (last column) reveals the power in the reconstructed microdata vis-à-vis the statistical baselines. There is only one substantive analytic difference between the COMRCL and $\text{CEF}_{atkr}$ results. Because we use the CEF itself to determine a nonmodal person, the data-defined populations in the COMRCL nonmodal and modal columns are a strict subset of the data-defined population in the all persons column: only the 106,300,000 records in the top left cell of Table 3 can be used to contrast the vulnerable (nonmodal) and nonvulnerable (modal) populations, whereas all 286,700,000 data-defined COMRCL records in the CEF universe can be used for the all persons column. All data-defined persons in the CEF can be studied in the $\text{CEF}_{atkr}$ results. For both attacker models, the denominators in the nonmodal and modal putative and confirmed reidentification rates from Equations 11 and 12 are the actual at-risk data-defined subpopulations (nonmodal and modal, respectively). The reidentification precision rate defined in Equation 13 is unaffected by restricting the domain of COMRCL. An external attacker using either COMRCL or higher quality data would still have to make inferences directly from the reconstructed microdata or form independent estimates of the race and ethnicity of the census block to estimate the putative or confirmed reidentification rate. But the external attacker’s estimate of the precision rate would still depend only on independent field work using a labeled sample of the attacker’s putative reidentification (see Section 5.2).
Figure 3. Comparison of Putative Reidentification Rates, Confirmed Reidentification Rates, and Reidentification Precision Rates for COMRCL by Census Block Size
Notes: The denominator used in the first column is the COMRCL Population column in Table LABEL:tab:reid_blockpop, which totals to 286,700,000. The denominator in the second column is the COMRCL population in Table LABEL:tab:reid_blockpop_nm, and the denominator in the third column is the COMRCL population in Table LABEL:tab:reid_blockpop_m. The denominators in the second and third columns sum to 106,300,000 because only that subset of COMRCL records can be classified as modal and nonmodal from the CEF. See Table 3.
@@
Notes: The denominator used in the first column is the $\text{CEF}_{atkr}$ Population column in Table LABEL:tab:reid_blockpop, which totals to 276,000,000. The denominator in the second column is the $\text{CEF}_{atkr}$ population in Table LABEL:tab:reid_blockpop_nm, and the denominator in the third column is the $\text{CEF}_{atkr}$ population in Table LABEL:tab:reid_blockpop_m. The denominators in the second and third columns sum to 276,000,000 because all records in $\text{CEF}_{atkr}$ can be classified as modal and nonmodal from the CEF. See Table 3.
For comparing and contrasting the nonmodal and modal persons results, the interpretation does not depend on the attacker database, so we characterize them generically. In these figures, nonmodal data-defined persons are the vulnerable population. Except for blocks with populations of 1-9 persons, CEF, HDF, $\text{rHDF}_{b,t}$, $\text{rHDF}_{b}$, MDG, and PRG putatively reidentify nonmodal persons at essentially the same rates because SF1 Tables P12 and 14 contain all the information used for putative reidentifications. It is the confirmation and precision rates that reveal the contrast with the modal person subpopulation. In blocks with 1-9 persons, the HDF, $\text{rHDF}_{b,t}$, and $\text{rHDF}_{b}$ have essentially identical confirmation and precision rates, whereas the MDG baseline is nearly zero for both confirmation and precision rates, and the PRG baseline is greater than the MDG but substantially less than either $\text{rHDF}_{b,t}$ or $\text{rHDF}_{b}$. As the population in the census block increases, the confirmation and precision rates of $\text{rHDF}_{b,t}$ and $\text{rHDF}_{b}$ decline but remain high, much closer to the rates for CEF and HDF than MDG (always essentially zero) and approaching proportional guessing only for the most populous census blocks. Even though there are 125 possible nonmodal {race, ethnicity} combinations, the reidentification precision rate of the reconstructed data is greater than 50% for 23,765,000 persons living in census blocks with populations of 1-99 persons (more than $1/3$ of all nonmodal data-defined persons) compared to essentially zero for the modal guesser. For these 23,765,000 persons, the smallest gain in the reidentification precision rate relative to the proportional guesser is 32.6 percentage points. [^34] Supplemental text Table 13, Census Block Size 50-99, COMRCL attacker, $\text{rHDF}_{b}$ reconstruction. This enormous gain in accuracy occurs only because the nonmodal person’s data were used in the SF1 tabulations. When the vulnerable population is defined as persons who differ from the predominant characteristics in small neighborhoods, in this case nonmodal {race, ethnicity} persons in lower-population census blocks, the statistical baselines either fail completely (MDG) or fair no better than chance (PRG), but the reconstructed microdata are correct at rates that approach the correctness of the confidential data themselves as the neighborhood population decreases.
@@ 7.3. Results for the most vulnerable populations
We now discuss the results for the most tightly constrained definition of the vulnerable population—nonmodal data-defined persons living in blocks with zero solution variability whose reconstructed records are population uniques for {block, sex, agebin}. Panel A of Table LABEL:tab:nm_solvar shows the results for nonmodal persons living in zero solution variability blocks and Panel B shows the results for the subset of these persons who are population unique on {block, sex, agebin}. Recalling that population uniqueness on {block, sex, agebin} implies population uniqueness on {block, sex, age} and that zero solution variability implies that the record in the reconstructed data is provably the same as the record in HDF without referencing any confidential data, Table LABEL:tab:nm_solvar Panel B establishes that the reidentification precision rates of $\text{rHDF}_{b,t}$ and HDF are essentially identical 95.4 versus 96.4 in COMRCL and 95.2 versus 95.6 in $\text{CEF}_{atkr}$. These reidentification precisions are far in excess of the modal guesser’s (3.3 for COMRCL and 2.5 for $\text{CEF}_{atkr}$) or the proportional guesser’s (20.9 and 20.2, resp.). Thus, the correct {race, ethnicity} has been exposed for more than 19 of 20 persons in the population of nonmodal, 0-$solvar$, {block, sex, agebin} uniques—essentially the same exposure as publishing the HDF itself, and only five percentage points worse than publishing the CEF itself. Statistical baselines are either almost always wrong or wrong for 4 of 5 persons in this vulnerable population. These results definitively show that the published SF1 tables result in confidentiality breaches since the high statistical precision of the reconstructed microdata is only possible because the target person’s data were used in the SF1 tabulations. If HDF was unsafe to release under the 2010 Census disclosure limitation framework, then the Summary File 1 statistics in Table 2 were also unsafe to release.
@@
Algorithm 3) (×$\times 10^{3}$) (×$\times 10^{3}$) (×$\times 10^{3}$) (%) (×$\times 10^{3}$) (×$\times 10^{3}$) (×$\times 10^{3}$) (%)
@@
requirements. The rows labeled MDF and $\text{rMDF}_{b,t}$ (light gray highlight) are the
analogs of $\text{rHDF}_{b,t}$ and HDF, respectively, when the 2020 Census DAS is applied
to the 2010 CEF. The rows labeled $\text{rSWAPLo}_{b,t}$ and $\text{rSWAPHi}_{b,t}$ (medium gray
@@
highlight) are the analogs of $\text{rHDF}_{b,t}$ when alternative swapping parameters are
used (see discussion in Section 9). Population for attacker COMRCL is data-defined
@@
nonmodal person records in COMRCL that match CEF on the feature set {pik, block}.
@@
Population for attacker $\text{CEF}_{atkr}$ is all data-defined nonmodal persons in CEF.
8. The 2010 Census Statistical Disclosure Limitation Framework
Did Not Meet the Census Bureau’s Own Standards
@@ 8. The 2010 Census Statistical Disclosure Limitation Framework
As indicated by the 95% binned-age agreement displayed in Table LABEL:tab:reconstruction_by_blocksize, it is possible to create an extremely accurate reconstruction of the underlying HDF microdata from publicly available sources, effectively undoing any confidentiality protection resulting from releasing tabular summaries instead of releasing the same features as a microdata extract of the universe of HDF person records. There is no reconstruction solution variability for $70\%$ of populated blocks, meaning that the reconstructed records in $\text{rHDF}_{b,t}$ were the exact image of the records in HDF for the feature set {block, sex, agebin, race, ethnicity}. For these blocks, housing-unit swapping and synthetic data for the group quarters population are the only sources of uncertainty about their correctness relative to the CEF. These SDL measures do affect the accuracy of the reconstruction relative to the CEF, especially in the blocks with the smallest population sizes (74% agreement in blocks with population size of 0-9), but the accuracy is still very high overall (and could be higher for attackers who attempt to undo the swapping). When the reconstructed data are linked to an external dataset, an attacker can learn extra characteristics about members of their dataset that they previously did not know, and could not learn statistically, with very high accuracy for select publicly identifiable subgroups. The attacker’s successful inference depends on several factors. First, the more accurately the external dataset matches the linking variables {block, sex, agebin}, the more records the attacker will successfully link and the better the inference about the value of the other census responses will be. In our experiment, only about $40\%$ of the records in our lower quality commercial data were correct matches to the CEF on these linking variables. Consequently, even though we could make a large number of putative matches, many were not confirmed—not because of incorrect inference on the unknown characteristics {race, ethnicity}, but because the characteristics on the external dataset were inaccurate quasi-identifiers. In our experiments using a much more accurate external database (either the $\text{CEF}_{atkr}$ or the subset of records in the commercial database that matched the CEF) the confirmation and precision rates were much higher. As the accuracy of the external data improves, the success of our attack improves, indicating that any conclusions based on the relatively lower quality of commercial name and address databases in 2010 should be tempered by the improved quality of such data by 2020.
@@
Our reconstruction-abetted reidentification attack was much more accurate in successfully inferring the $\{race,ethnicity\}$ of records with a value that was different than the most common (modal) $\{race,ethnicity\}$ in the block in comparison with the baseline inferential methods of guessing the modal {race, ethnicity} and guessing the {race, ethnicity} with probability proportional to publicly released counts. It also vastly outperformed the baseline methods when making inferences about records that were unique and had nonmodal $\{race,ethnicity\}$, facts that the attacker could infer exactly in blocks with zero solution variability without consulting any confidential data. Therefore, while it is possible to make accurate inferences with these baseline methods when statistical information is used, the reconstruction-abetted reidentification attack exposes the vulnerability of the confidentiality protections by permitting accurate inferences on persons with statistically uncommon and rare characteristics. The accurate inference of the race and ethnicity of a person who is a population unique is only possible because that person’s record was available in $\text{rHDF}_{b,t}$. Such an inference is nonstatistical using the framework we laid out previously (see Section 5). Record-level data swapping marginally reduced the number of records in the $\text{rHDF}_{b,t}$ at risk for a putative match, but it had no discernible effect on the precision of nonstatistical inferences for the vast majority of records.
@@
The conclusion from this analysis is that the disclosure avoidance system applied to the 2010 Census did not meet the requirements for approval set forth by the Census Bureau’s Data Stewardship Executive Policy (DSEP) Committee for 2010 Census publications and, in light of current algorithmic and computational capabilities, it is not viable for future use. The results of the this experiment motivated the Census Bureau to change the SDL methods used for the 2020 Census (abowd2018us). Instead of record swapping, synthetic group quarters data, and aggregation, the Bureau used differentially private mechanisms for all tabulations, microdata, and other statistical products computed from the 2020 Census (Abowd20222020). Had the Census Bureau continued to use the 2010 technology for 2020, it would have had to use the same rules for microdata as for tabular data, likely resulting in a much higher swap rate and a large amount of suppression of block-level data.
9. The 2020 Census Statistical Disclosure Limitation System
Addressed the Failures of the 2010 System
@@ 9. The 2020 Census Statistical Disclosure Limitation System
Our final discussion addresses whether any SDL method can simultaneously defend against reidentification attacks and other nonstatistical uses while allowing publication of data that are fit for use in the primary applications. christ:radway:bellovin:2020 demonstrated that the Census Bureau’s differentially private disclosure avoidance system could provide fitness-for-use in the redistricting application comparable to swapping but with superior confidentiality protection. They also noted that “[s]wapping places a disproportionate privacy burden on minority groups,” while differential privacy protections apply to all sub-populations. These results, which we confirm in this paper, are in contrast to the erroneous conclusions drawn by kenny:et:al:2021 who failed to distinguish between generalizable and confidentiality-violating inferences (jarmin:etal:2023). wright:irimata:2021 demonstrated that the 2020 Census Redistricting Data (P.L. 94-171) Summary File was fit for use in redistricting and scrutiny under Section 2 of the 1965 Voting Rights Act (vra).
@@
On October 20, 2022, DSEP approved the final production settings for the 2020 Census disclosure avoidance system as applied to the Demographic and Housing Characteristics (DHC) File, the successor to the 2010 Summary File 1 used in this paper. DHC was produced using an extension of the algorithms described in Abowd20222020 that includes the 2020 Census version of the tables used in this paper (cumings-menon:etal:2023). The statistical summaries presented for the DSEP decision included an analysis of the effectiveness of the same reconstruction-abetted reidentification attack studied here based on the 2010 Census. The analysis compared the confidentiality protections from the original SF1 tables shown in Tables 2, 3 and 4 with the protection afforded by the differentially private disclosure avoidance system, when executed at the DHC production settings applied to the 2010 Census data, and the protection afforded by housing unit-level data swapping at rates of 5% and $50\%$, i.e., swapping 5% or half of all housing units to a new location in a different block, tract, or elsewhere in the state (for details, see Appendix C).
@@
Table LABEL:tab:agree_blockpop_mdf shows the agreement rates overall and by census block size between the 2010 CEF and
@@
  • $\text{rMDF}_{b,t}$—reconstructed Microdata Detail File (MDF), (Abowd20222020), the production version output of the differentially private DAS for the DHC when applied to 2010 Census data and processed into the equivalents of the Table 2 statistics,
@@
The $\text{rMDF}_{b,t}$, MDF, and $\text{rSWAPHi}_{b,t}$ reduce the agreement with the 2010 CEF substantially, with the $\text{rMDF}_{b,t}$ and MDF reducing agreement more than $\text{rSWAPHi}_{b,t}$. The $\text{rSWAPLo}_{b,t}$ agrees with the CEF at essentially the same rate as $\text{rHDF}_{b,t}$ except for blocks with populations of 1-9 persons, where HDF, $\text{rHDF}_{b,t}$, and $\text{rHDF}_{b}$ all agree with the CEF for 74.0% of records using binned age while $\text{rSWAPLo}_{b,t}$ agrees with 94.9% of these records. The experimental swap methodology did not single-out low-population blocks for relatively greater swapping.
Table 11. Selected Reconstruction Agreement Statistics with Comparisons to Output from the 2020 DAS and Specially Swapped Versions of the CEF Using the 2010 Census as Input by Census Block Size
@@
Data ($L-R$ in Block Population Agreement (×$\times 10^{3}$) Agreement (%)
Algorithm 2) Size (×$\times 10^{3}$) Exact Age Binned Age Exact Age Binned Age
@@
$\text{rSWAPLo}_{b,t}$ (5%)-CEF All 308,746 144,500 281,200 46.8 91.1
$\text{rSWAPHi}_{b,t}$ (50%)-CEF All 308,746 100,800 193,200 32.7 62.6
@@
$\text{rSWAPLo}_{b,t}$ (5%)-CEF 1-9 8,070 3,278 7,660 40.6 94.9
$\text{rSWAPHi}_{b,t}$ (50%)-CEF 1-9 8,070 1,803 4,280 22.3 53.0
@@
$\text{rSWAPLo}_{b,t}$ (5%)-CEF 10-49 67,598 30,320 63,370 44.9 93.8
$\text{rSWAPHi}_{b,t}$ (50%)-CEF 10-49 67,598 18,110 38,330 26.8 56.7
@@
$\text{rSWAPLo}_{b,t}$ (5%)-CEF 50-99 69,073 31,190 63,300 45.2 91.6
$\text{rSWAPHi}_{b,t}$ (50%)-CEF 50-99 69,073 20,200 41,180 29.2 59.6
@@
$\text{rSWAPLo}_{b,t}$ (5%)-CEF 100-249 80,021 36,310 71,880 45.4 89.8
$\text{rSWAPHi}_{b,t}$ (50%)-CEF 100-249 80,021 25,740 50,530 32.2 63.2
@@
$\text{rSWAPLo}_{b,t}$ (5%)-CEF 250-499 42,911 20,470 38,250 47.7 89.2
$\text{rSWAPHi}_{b,t}$ (50%)-CEF 250-499 42,911 15,830 29,030 36.9 67.7